When Beza's income is Birr 2000; she purchases 100Kg of a good. When her income increases to Birr 2400, she purchases 120Kg of a good. A) What is her income elasticity? B) What type of good is?


1
Expert's answer
2022-01-05T11:13:47-0500

Income elasticity of demand refers to the sensitivity of the quantity demanded for a certain good to a change in the real income of consumers who buy this good. The formula for calculating income elasticity of demand is the percent change in quantity demanded divided by the percent change in income.

E = ((120-100) / 120) / ((2400-2000)/2000) = 0.17 / 0.2 = 0.85

An inferior good has an Income Elasticity of Demand < 0.